# EcoSense: Energy-Efficient Intelligent Sensing for In-Shore Ship Detection through Edge-Cloud Collaboration

### v1

@@ 1. Introduction
Apart from algorithmic challenges, IMOD also suffers from difficulties in real-world deployments. Sensors for IMOD, such as coastal webcams, drones, and movable cameras, are deployed in vast numbers and in challenging environments along lengthy and narrow coastlines. Such an elevated complexity of the IMOD systems requires the awareness of platform and environmental constraints and optimized utilization of computational resources. In addition to the techniques of image compression (Zhang et al., 2023; Mao et al., 2022), previous efforts to tackle this challenge can generally be categorized into two approaches: centralized and distributed, as illustrated in Fig. 1. While both approaches alleviate the problem to some extent, they exhibit intrinsic drawbacks. In a centralized approach (Fig. 1(a)), the detection model is deployed on a central powerful server (i.e., cloud), and all sensors transmit data to the server for inference. Despite delivering outstanding performance, this approach escalates the volume of data transferred over the wireless channel, demanding high bandwidth and reliability. Furthermore, any anomaly occurring on the server can lead to a system-wide shutdown. Conversely, a distributed approach, as presented in Fig. 1(b) and explored in (Sharafaldeen et al., 2022), involves deploying the detection model in proximity to each sensor. In this configuration, the inference process occurs near the sensor, with only the results transmitted externally. While this approach conserves bandwidth and enhances resilience against potential system failures, it grapples with challenges related to battery life, execution time, and overall performance owing to limitations imposed by edge devices.
@@ 2.2. Object Detection
Object detection is one branch of computer vision widely applied in people’s lives, such as monitoring security (Joshi et al., 2012). The majority of CNN-based object detection models can generally be categorized into two classes: two-stage detectors (Girshick, 2015) and one-stage detectors (Redmon et al., 2016). Two-stage detectors such as Faster R-CNN (Ren et al., 2015) first suggest candidate object bounding boxes via a Region of Interest (RoI) pooling layer, and the second stage involves classification and refinement processes upon features extracted from bounding boxes. In contrast, one-stage detectors forego the region proposal and directly perform detection over a dense sampling of locations (Redmon et al., 2016). On the other hand, the emergence of Transformer architecture has demonstrated its superiority over conventional CNN methods in certain tasks (Liu et al., 2021b). Nevertheless, the deployment of these models on resource-constrained devices is hindered due to their substantial computational demands. Although efforts have been made to develop lightweight variants, such endeavors often lead to compromised performance (Han et al., 2020; Hu et al., 2018).
@@ 3. Method
Figure 2. Front-end model (a). Schematic of front-end model (b). Modified Ghost module. Φ()normal-$\Phi(\cdot)$ denotes the linear projection in the original Ghost module. Red arrows show the path of the Squeeze-Excitation operation, and blue arrows show Coordinate Attention ($F_{sq}(\cdot)$ denotes squeeze operation, $F_{ex}(\cdot)$ denotes excitation operation). $\odot$ denotes element-wise multiplication.
@@ 3.2. Back-end Model
Given the high similarity between categories, we leverage a robust Swin-Transformer (Liu et al., 2021b) combined with a Graph Convolutional Network (GCN) for improved classification of hard samples, inspired by the approach presented in (Jia et al., 2023). A schematic of our back-end model is depicted in Fig. 3. However, prior models often concentrated on localizing subtle discrepancies while neglecting the valuable information provided by the background. This information can indicate which features are unnecessary or potentially harmful for accurate classification. To address the limitations, we introduced a Background Suppression (BS) module aimed at mitigating unnecessary background interference while augmenting the representation of diverse features (Chou et al., 2023).

### v3

@@ 1. Introduction
Apart from algorithmic challenges, IMOD also suffers from difficulties in real-world deployments. Sensors for IMOD, such as coastal webcams, drones, and movable cameras, are deployed in vast numbers and in challenging environments along lengthy and narrow coastlines. Such an elevated complexity of the IMOD systems requires the awareness of platform and environmental constraints and optimized utilization of computational resources. In addition to the techniques of image compression (Zhang et al., 2023; Mao et al., 2022) and memory management for edge devices (Wen et al., 2020, 2021b, 2022, 2021a), previous efforts to tackle this challenge can generally be categorized into two approaches: centralized and distributed, as illustrated in Fig. 1. While both approaches alleviate the problem to some extent, they exhibit intrinsic drawbacks. In a centralized approach (Fig. 1(a)), the detection model is deployed on a central powerful server (i.e., cloud), and all sensors transmit data to the server for inference. Despite delivering outstanding performance, this approach escalates the volume of data transferred over the wireless channel, demanding high bandwidth and reliability. Furthermore, any anomaly occurring on the server can lead to a system-wide shutdown. Conversely, a distributed approach, as presented in Fig. 1(b) and explored in (Sharafaldeen et al., 2022; Liu et al., 2024), involves deploying the detection model in proximity to each sensor. In this configuration, the inference process occurs near the sensor, with only the results transmitted externally. While this approach conserves bandwidth and enhances resilience against potential system failures, it grapples with challenges related to battery life, execution time, and overall performance owing to limitations imposed by edge devices.
@@ 2.2. Object Detection
Object detection is one branch of computer vision widely applied in people’s lives, such as monitoring security (Joshi et al., 2012). The majority of CNN-based object detection models can generally be categorized into two classes: two-stage detectors (Girshick, 2015) and one-stage detectors (Redmon et al., 2016). Two-stage detectors such as Faster R-CNN (Ren et al., 2015) first suggest candidate object bounding boxes via a Region of Interest (RoI) pooling layer, and the second stage involves classification and refinement processes upon features extracted from bounding boxes. In contrast, one-stage detectors forego the region proposal and directly perform detection over a dense sampling of locations (Redmon et al., 2016). On the other hand, the emergence of Transformer architecture has demonstrated its superiority over conventional CNN methods in certain tasks (Liu et al., 2021b). (Yu et al., 2024) has shown the effectiveness of combining Transformer models with T-SNE to build efficient detection models, highlighting the critical role of these models in improving detection accuracy and efficiency by capturing complex data patterns. Nevertheless, the deployment of these models on resource-constrained devices is hindered due to their substantial computational demands. Although efforts have been made to develop lightweight variants, such endeavors often lead to compromised performance (Han et al., 2020; Hu et al., 2018).
@@ 3. Method
Figure 2. Front-end model (a). Schematic of front-end model (b). Modified Ghost module. $\Phi(\cdot)$ denotes the linear projection in the original Ghost module. Red arrows show the path of the Squeeze-Excitation operation, and blue arrows show Coordinate Attention ($F_{sq}(\cdot)$ denotes squeeze operation, $F_{ex}(\cdot)$ denotes excitation operation). $\odot$ denotes element-wise multiplication.
@@ 3.2. Back-end Model
Given the high similarity between categories, we leverage a robust Swin-Transformer (Liu et al., 2021b) combined with a Graph Convolutional Network (GCN) for improved classification of hard samples, inspired by the approach presented in (Jia et al., 2023). In addition, we follow (Zheng et al., 2024), which utilizes the Synthetic Minority Over-sampling Technique (SMOTE) combined with XGBoost to improve the robustness. A schematic of our back-end model is depicted in Fig. 3. However, prior models often concentrated on localizing subtle discrepancies while neglecting the valuable information provided by the background. This information can indicate which features are unnecessary or potentially harmful for accurate classification. To address the limitations, we introduced a Background Suppression (BS) module aimed at mitigating unnecessary background interference while augmenting the representation of diverse features (Chou et al., 2023).